construct a sampling distribution of the sample means such that the average number of mg of cholesterol in a cup of a certain brand of ice cream is 660mg and the standard deviation is 35mg


1
Expert's answer
2021-06-08T09:01:33-0400

Solution:

Given, "\mu=660\ mg, \sigma=35\ mg"

Now sampling distribution for sample size "n" :

"\mu_{\bar X}=\mu=660\ mg
\\\sigma_{\bar X}=\dfrac{\sigma}{\sqrt n}=\dfrac{35}{\sqrt n}"